Suppose that the economy is characterized by the following behavioral equations: C = 150 + 0.4Yd-6i, I = 70 + 0.2Y-i, G = 80, T = 15+0.25Y, Md = 0.2Y-15-75i, M=2800,P=20
1
Expert's answer
2019-11-05T09:55:26-0500

C = 150 + 0.4Yd-6i

I = 70 + 0.2Y-i

G = 80

T = 15+0.25Y

Md = 0.2Y-15-75i

M=2800

P=20

Yd = Y-T

Y = C+I+G = 150 + 0.4Yd-6i + 70 + 0.2Y-i +80 = 300 +0.4Yd -7i+0.2Y = 300 +0.4(Y-15-0.25Y)-7i+0.2Y = 300 +0.3Y -6 -7i+0.2Y

0.5Y = 294 -7i

Y = 588 - 14i

2 i = 70+0.2Y

Y = 588 - 980 - 1.4Y

Y = 163.3